A yeast manufacturer finds that the yeast will grow exponentially at a rate of 12.1% per hour. How many kilograms of yeast must the manufacturer start with to obtain 500 kg at the end of 8 h?
Round to the nearest kg.

Answer by Boreal(15235)

500=A(1+.121)^8
500=A*2.4937
A=500/2.4937
=200.5049 or 200 kg